OpenStudy (anonymous):

Find the slope of the line passing through (-3, 2) and (2, 4).

OpenStudy (luigi0210):

y2-y1/x2-x1

OpenStudy (luigi0210):

where y2=4, y1=2, x2=2 and x1=-3

OpenStudy (mathstudent55):

Slope is the difference in y values over (divided by) the difference in x values. Your y values are 2 and 4 Your x values are -3 and 2 Difference in y: 2 - 4 = -2 Difference in x: -3 - 2 = -5 Slope = -2 / (-5) = 2/5

OpenStudy (mathstudent55):

Example: Find the slope of the line passing through (-8, 10) and (-8, -7). slope = (10 - (-7) ) / (-8 - (-8) ) = 17 / 0 Since this slope results in the division by zero, the slope is undefined.

OpenStudy (mathstudent55):

Another example: Find the slope of the line passing through (0, -5) and (-4, 7). slope = (7 - (-5) ) / (-4 - 0) = (7 + 5) / (-3) = 12 / (-3) = -4